THE 4^{3}Δ_{g} STATE IN K_{2} - INVESTIGATING A POSSIBLE GATEWAY TO CORE NON-PENETRATING RYDBERG STATES

\underline{JENNY M. MAGNES}, Department of Physics, United States Military Academy, West Point, New York 10996, USA; A. MARJATTA LYYRA, ERGIN AHMED, Department of Physics, Temple University, Philadelphia, PA 19122; YAOMING LIU, Department of Chemistry, University of Illinois at Chicago, Chicago, IL 60607; LI LI, Tsinghua University,Key Lab Atom and Molecular Nanoscience, Beijing 100084, Peoples Republic of China; SYLVIE MAGNIER, I.U.M.F. de Bretagne, 153 Rue Saint Malo, CS 54310 F-35042 Rennes Cedex, France, and, Laboratoire de Physique des Atomes, Lasers, Molécules et Surfaces (PALMES), CNRS et Université Rennes I (UMR6627), Campus de Beaulieu, Bât. 11B, F-35043 Rennes Cedex, France; MONIQUE FRÉCON, Laboratoire de Spectrométrie Ionique et Moleculaire (L.A.S.I.M), CNRS et Université Lyon (UMR5579), Campus de la Doua, Bât. Alfred Kastler, 43 Bd du 11 Novembre, F-69622 Villeurbanne Cedex, France.

Core non-penetrating Rydberg states can give useful information on the electronic structure of the ion core; however, core non-penetrating states are difficult to observe since these states hardly penetrate the more accessible ion core and the electronic angular momentum quantum number, l, is large, for the core non-penetrating states thus the transition dipole moment to the core non-penetrating states is small. The core penetrating 4^{3}Δ_{g} state (atomic limit: 4s+5d) and the core non-penetrating 3^{3}Δ_{g} state (atomic limit: 4s+4f) perturb each other since they have the same symmetry and overlapping energy states thus creating the possibility of a gateway to other core non-penetrating states.